[ Japanese | English ]

過去のお知らせ

アンテナ能率 (2015 - 2016)に低周波数帯の値を追記しました (2016-06-12)
アンテナ能率 (2015 - 2016)を公開しました (2016-05-17)
ポジションスイッチ観測におけるバグが見つかりました。 2012年から2016年の間に該当する観測を行った方はリンク先より確認をお願いいたします(日本語はリンク先下部にあります)。(2016-05-02)
連続波バックエンドのステータスを更新しました。#11は使用を避けてください。 (2016-03-21)
受信機設定のサンプルファイルページを観測スクリプト作成ツールObstable ZのDevice Tabページからリンクしました。(2016-01-19)
観測マニュアルをアップデートしました。 (2015-12-03)
バグ修正版Time Estimator for OTFをリリースしました。
OTF観測の提案を考えていらっしゃる方はこちらを使って観測時間の再見積りをしていただくようお願いいたします。 (2015-11-30)
Time Estimatorのバグ修正のため、共同利用後期の締め切りを12月1日から12月7日へ延期しました。
バグ修正版のリリースは11月30日を予定しております。(2015-11-27)
観測スクリプト作成ツールObstable Zのページを改定しました。(2015-11-17)
プロポーザル採択後、観測までの準備ガイドのページを公開しました。(2015-11-17)
2015-2016シーズン前期のプロポーザルの要求時間をLST帯ごとにまとめた図をプロポーザル情報のページに公開しました (2015-11-13)
プロポーザル提出に関してよくある質問をFAQとしてまとめました (2015-11-02)
最新のアンテナのスカイラインを観測案作成ガイドページに記載しました(2015-09-04)
45m電波望遠鏡共同利用公募のお知らせ(2015-08-03)→プロポーザル情報(Call for Proposals)
アンテナ能率 (2014 - 2015)を公開しました (2015-05-26)
helpdeskを設置しました (2015-03-06)
教育支援枠のプロポーザルを、12月10日締切で追加募集いたします (2014-10-22)
新マルチビーム受信機FORESTの情報 (2014-08-14)
45m電波望遠鏡共同利用公募のお知らせ (2014-07-31) →プロポーザル情報 (Call for Proposals)
標準天体スペクトルリストを更新しました。(2014-04-25)
野辺山観測所データ解析サーバーご利用の皆様へ(野辺山観測所データ解析サーバ運用終了のお知らせ) (2013-05-16)
新分光器SAM45のステータス
AOSのステータス (2012-03-02)
AOS-W3が故障しています (2014-02-04)
アンテナ能率を公開しました (2014-04-17)
S80受信機は故障のため、現在使用できません。 (2014-03-03)
なお、S80受信機の修理及び性能確認を行うため、同じデュワーに入っているS100受信機もご使用いただけなくなります。ご迷惑をおかけします。現在調査がやや長引いております。
  → 修理及び調査を行い正常に機能することを確認し、運用を再開しました。(2014-04-02)
新システムのADコンバーター(PANDA)が一台新型になり、IF 13, 14に使われています(PANDAによるスプリアスが見られますが、このPANDAについてはスプリアスが出なくなりました) (2013-12-24)
採択されたプロポーザル (2013-10-16)
45m電波望遠鏡共同利用公募のお知らせ (2013-07-24) →プロポーザル情報 (Call for Proposals)
電波シーイングモニタが見れるようになりました。(観測所内からのみアクセスできます) (2013-05-20)
アンテナ能率を公開しました (2013-04-12)
現在、BEARSのビームの1つであるch3-5は、第一局部発振器が108 GHz以上で性能が出ません。観測計画を立てる際にはご注意ください (このビームでの12COの観測は難しいですが、13COをUSBで観測するのは可能です)。(2013-02-27)
3 mm帯の2SB受信機T100の運用を終了しました。今後は、同等の機能を持つTZ1受信機をお使いください。TZ1はSAM45分光器と接続することができます。現在、関連のweb pageを更新中です。TZ1の観測テーブル作成には、ソフトとしてObstable Z (zobs)をご利用ください (2013-01-08)
Obstable Z (zobs)でのdevice tableの作成方法が変わりました。こちらのページをご参照ください。(2013-01-08)
新マルチビーム受信機FORESTの情報(2012-08-29)
CASA reduction manual (2011/12/16)
Spectra of Standard Sources is here (2011-01-14)
Java version of Newstar (reduction software for spectral lines) was released. In addition, a task (beta version) of Newstar "New Spectrum FITS" (to transform spectral files to FITS format) was also released.
For details click here. (2010-01-06)
How to start OBSTABLE for 45m is here.
The proposal cover sheet has been upgraded. Please use ver 4.10.(2008-10-30)
Proposal information for open-use observation in 2009 season is now open.(2008-09-03)
The deadline of General Program has been postponed to 11 November 2008.
The LO of H30 receiver does not work stably in some frequency. Please ask us for details. (2008-05-20)
IF1 attenuator is broken (May 6-13, 2008)
The 1dB-step attenuator of IF #1 is broken. Please do not use IF 1.
Export version of NOSTAR
Export version of OTF reduction software is released. (Mar. 31, 2008)
Antenna efficiencies for 2007-2008 season
The antenna efficiencies in the 2007-2008 season is presented. (Mar. 13, 2008)
A bandpass filter of IF2 cannot be used. (Feb. 25-26, 2008)
The 5.1--5.6 GHz bandpass filter of the IF 2 was broken, and seems to be "spontaneously" repaired.
AOS-H7 cannot be used. (Dec. 29, 2007-Feb. 5, 2008)
The LO attenuator of the AOS-H6 is broken.
AOS-H6 cannot be used. (Jan. 3-8, 2008)
The step attenuator of the AOS-H6 sometimes does not function.
Export (beta) version of NOSTAR
Public beta release of Export version of OTF reduction software. (2007-11-30)
BEARS star formation project
Data from BEARS star formation project is open. (2007-08-28)
Status report
Status Report for 2007-2008 season is now open.(2007-06-15)
Proposal information
Proposal information for open-use observation in 2007-2008 season is now open. (2007-06-06)
IF7 attenuator is broken (Apr. 27-May 1, 2007)
The 1dB-step attenuator of IF #7 is broken. Please do not use IF 7.
Beam efficiencies for 2006-2007 season
Beam efficiencies are released. (2007-04-25)
BEARS scaling factors for 2006-2007 season
BEARS calibration data (scaling factors) are released. (2007-04-19)
BEARS (CH15) cannot be used. (Mar. 19-21, 2007)
Due to a snapping of an IF cable, CH15 of BEARS cannot be used since Mar. 19. And a trouble of the refrigerator which occurred succeedingly makes all the 25 beams unavailable (Mar. 20). Both troubles are fixed by Mar. 21.
The AOS-H7 cannot be used (Mar. 1-13, 2007)
The 4th local oscillator for AOS-H7 is broken.
IF2 attenuator is broken (Feb. 22-27, 2007)
The 1dB-step attenuator of IF #2 is broken. Please do not use IF 2.
Trouble on AC-03 (Feb. 13-14, 2007)
Due to a trouble on a digital spectrometer A03, CH13 of BEARS cannot be used.
We have renewed the Web site. (Feb. 7, 2007)
Old page (Japanese)
Status of AOS-W1 (Feb. 6-14, 2007)
The band characteristics of AOS-W1 was found to be bad. Please do not assign important lines only in this spectrometer.
--> It turned out to be OK.
Operation resumed
The 45 m telescope was temporary repaired, and the operation (common use observations) has started again. For details see the top page of this observatory. (Feb. 5, 2007)
Operation Interrupted (Jan.-Feb. 5, 2007)
The operation of the 45 m telescope is now stopped due to the trouble of the driving system. For details see the top page of this observatory. (Jan. 24, 2007)
Schedule of open use observations in 2006-2007 season
Open use will start from early Jan. 2007. Now we are preparing for it. (Dec. 24, 2006)
[ collimator: done | subref: done | pointing: done | beam squint: done | beam efficiencies: almost done | etc. ]
We now open the OTF observing mode.
OTF (On The Fly) mode is available for open use.
Change of 45-m manager
N. Kuno has became the manager of the telescope in succession to K. Sunada. (Nov. 2006)
The frequency converter of AOS-H3 is now broken.
Unfortunitely AOS-H3 cannot be used. (March 28 - April 4, 2006)
The H40 receiver was replaced.
The 1st IF and the side band were changed (9.8 ---> 6.0 GHz, LSB ---> USB). You can get a sample file from the window of the device table.
Present status of AOS (Dec. 2004-)
Present status of BEARS
At present a channel 25 cannot be used due to the breaking of the wire (Dec. 31, 2004-Jan. 27, 2005).
The intensity calibration system of the old port (old MMC) is broken (Jan. 8-19, 2005)
Please use LOAD1 (LD1) instead of the old MMC for the observations with the H40, H32/28, and H22 receivers.
For this coming observational season, we are measuring the telescope performance and adjusting the basic parameters (subreflector, pointing, beam squint, efficiency, etc.).
The adjustment of the subreflector position and the revision of pointing parameters have already finished.
The operation of the 45m radiotelescope in this season finished. (31 May 2004-)
On 27 May the operation in this observational season finished. In addition to the usual maintenance works, we are going to replace the control system of the telescope in this summer.
Trouble of AOS-W4 (13 April 2004-)
The baseline of AOS-W4 may be bad. Please do not assign important lines only in this spectrometer.
The revisions of OBSTABLE (2 Dec. 2003-)
The use of the digital backend for the single beam receivers (2 Dec. 2003-)
From this observing season, it is possible to connect the digital backend to the single beam receivers. The manual will be available soon. For details please ask to Dr. Sunada.
Present status of AOS (Dec. 2003-)
Trouble of AOS-W4 (9-13 April 2004)
At present fringe noise appears at half of the channel ranges of AOS-W4. Please do not assign important lines only in this spectrometer.
Trouble of AOS-H8 (2 Dec. 2003 - 9 March 2004)
At present the final local oscillator for AOS-H8 is broken, and consequently, H8 cannot be used.
Trouble of an attenuator at IF7 6-2 GHz converter (8-9 Dec. 2003)
Please do not use the IF7, or use it with even number (dB) of attenuation in the device table (ask to the assistant). In this case, please do not assign important lines only in the IF7.
Now the 45m telescope is not operated for the maintenance. (deleted 28 Nov. 2003)
Problem of the band selection board (1 April 2003 -)
The S100 receiver cannot be connected to the IF 7 line. Please do not assign the IF 7 line for the S100 receiver.
PLL problem of the S100 receiver (18 March 2003 -)
The PLL of S100 does not work well at the local frequency of 87.925 GHz. Please avoid to use this local frequency.
Tuning of the S100/S80 receivers (Dec. 2002 -)
Almost complete tuning can be done by PC at the control room. In this tuning process bias voltage is not adjusted. If this adjustment is necessary to obtain maximum performance, please ask to the assistant.
Present status of BEARS
Information on the local oscillators of the H22 receiver
Please note that the tuning ranges of the local (Gunn) oscillators of H22 is not completely continuous. The tuning ranges are 14.75-15.25 GHz, 16.7-17.25 GHz, and 18.7-19.3 GHz. Their will be no big problem, because the bandwidth of H22 is rather wide (2 GHz).
Present status of AOS (Dec. 2002-
Connection of cables for continuum observations (NOBA and other receivers)(Dec. 2002-)
This connection became automatic by using a switch box. Please click 'start tuning' in the main console to make this connection effective before continuum observations.
Low performance of the S100 receiver (ca. 11 March 2003 -) (deleted 1 April 2003)
The performance of S100 is unfortunitely slightly low at present. Roughly speaking it is better to use S80 instead of S100 for observations below 100 GHz. For observations above 100 GHz S100 seems better than S80. For details please ask the people maintaining the receivers.
Trouble of IF6 (19 March 2003 -) (deleted 25 March 2003)
The system temperature of AOSs connected to IF6 is now abnormally low. Please avoid to use IF6.
Trouble of an attenuator (14 March 2003-) (deleted 17 March 2003)
An attenuator of IF6 at 6 to 2 GHz converter is now broken. Please do not use the IF6 line.
Present status of BEARS (deleted 18 Feb. 2003)
Trouble of an attenuator at IF8 6-2 GHz converter (7-17 Feb. 2003)
In case of odd number (dB) of attenuation, signal does not pass. Please do not use the IF8, or use it with even number (dB) of attenuation (ask to the assistant).

2016-06-21 Update